30 January 2010: Active Optics tests

05:10 Open. Cloudy skies, but humidity dropped below threshold

05:13 Slew to ACT0370, pointing star. Some bumps with observing in only blue... maybe because I forgot to untick red channel and mirror (though Red was not powered up), or because PSF Right was not in simulate? Too many commands - restarted blue

05:29 first dofpia, /first, /blueonly, bluelimit=600 - no good pupils, try with /x2. Still no luck - brightest star has about 39,000 cnts, background 32,000. Going to M35 instead.

05:30 Note that az/el numbers on User Interface are both 0.0000. But, during the slew, reasonable values flashed up and then 0.0000 came back.

05:55 dofpia seems to be selecting some of the fainter, more blended pupils and ignoring the brighter more isolated ones. See 054953 - 055410. Stop with "S" --- but last image, 05558, is better.

06:00 move N by 1arcmin... run dofpia again, without /x2.

06:50 So far no success in collimating --- the pupils which are 'found' are not the best ones and the fits are not good.

07:00 Back to a pointing star, dofpia converged - to relaxed criteria because of poor seeing (2.38" it says). Observe ACT0331 to check/correct pointing.

07:30 Again, 3 pupils found but which do not look like the best pupils. See, e.g. lbcb.20100130.072852.fits

09:10 Drove in Z22 = -200nm --- as we had been consistently seeing a ring.

  3724    368   1087     91    808  -1000   -100    y  1.24  1.67                      /newdata/lbcb.20100130.091758.fits  V-BESSEL
 -4805   -805   -595    -85   -143   1000    100    y  1.19  2.02                      /newdata/lbcb.20100130.092001.fits  V-BESSEL

09:30 2nd image I noticed with lots of hot pixels. Number is 092933.

09:38 Converged pupils have positive Z22; we had found the zeropoint of 2.3 was not correct in October and did some initial tests with 1.5 (but we had not tuned focus). We'll change zeropoint to 1.5 now and see how this affects dofpia convergence.

10:20 Moving to NGC2841 to test the Active Optics on tech chips... But note, I left Zeropoint = 1.5. Taking guide chip images -

10:47 to a new part of sky, galactic coords 200, 40 where I thought there would be more stars. There must be clouds - upped the integration time to 2x. Clouds in this part of sky - illuminated by moon and saturated.

11:10 Cleared Active Optics (large tip/tilt had built up) and running dofpia again. Lot of coma --- Gave +2000 Z7 but this didn't help 112101 Gave -4000 Z7 and see more coma (rotator angle was -60 and bright area opposite spider). -112451. Gave +8000 Z7; this caused Y limit to be hit. Wayne had to do coma-free pointing correction.

11:50 dofpia not converging well - stopping calibration tests and going to test whether the active optics program works. On field at 14h, 25d, elevation=76deg. Not well collimated --- just want to see how program runs. 12:15... running set of 60sec integrations to activate tech chips

12:20 We note in the lbcfpiaBlueLog a sequence where there is some oscillation of Z4 and Z11 in phase with Z22. This may be helped if there would be a damping factor for Z22... for low Z22 (within noise), it has less effect - is not sent and has less effect on Z4 and Z11.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   -38  -1843   -697    -53    -15    344     85    y  1.24  0.62                      /newdata/lbcb.20100130.115927.fits  V-BESSEL
  -363   -431   -488   -212     14    191     66    y  1.05  0.30                      /newdata/lbcb.20100130.120047.fits  V-BESSEL
  1197    190    335    275     39   -426    -60    y  1.07  0.57                      /newdata/lbcb.20100130.120207.fits  V-BESSEL
 -1100    187    183    -82    109    282     62    y  1.14  0.49                      /newdata/lbcb.20100130.120328.fits  V-BESSEL
   808     96   -411     99   -207   -193     -4    y  1.07  0.39                      /newdata/lbcb.20100130.120447.fits  V-BESSEL
  -315    564    152   -382     11    -38    -18    y  1.09  0.28                      /newdata/lbcb.20100130.120608.fits  V-BESSEL
  -213   -819    478    131     11     53     35    y  1.08  0.31                      /newdata/lbcb.20100130.120731.fits  V-BESSEL
  1016   -314   -574    183    105   -515    -76    y  1.02  0.56                      /newdata/lbcb.20100130.120852.fits  V-BESSEL
 -1225      7    767   -197     29    379     22    y  0.98  0.60                      /newdata/lbcb.20100130.121014.fits  V-BESSEL
   552   -687   -734   -304    274      0      0    y  2.01  0.42                      /newdata/lbcb.20100130.121133.fits  V-BESSEL
  1906      0      0      0      0  -1003      0    y  0.00  0.00                     dofpia_sph3_adjustment             

12:24 We saw no change in Z4 during the OB at 12:15 --- forgot to stop/restart LBC. Doing that now. 12:30 or so --- we ran the OB again, but see no changes in Z4 --- the tech chip images do not have stars or very many stars. We are going to a field closer to the galactic plane - 16h 00d

13:05 Telescope stopped tracking during 1259 exposure. Azimuth is stuck.

-- OlgaKuhn - 30 Jan 2010
Topic revision: r4 - 31 Jan 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback